# parskip=half

### v2

@@ 2.2 Conditionally independent censoring
An alternative hypothesis in survival analysis that relaxes the assumption of independent censoring is conditionally independent censoring, also referred sometimes as informative censoring. It allows to model more realistic censoring processes, in particular in situations where there are reasons to believe that $C$ may be dependent on $A$ and $X$ (for instance, if a patient is more likely to leave the study when treated because of side-effects of the treatment).
@@
(Conditionally Independent Censoring)
$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X,A$
@@
(Positivity / Overlap for Censoring)
There exists $\varepsilon>0$ such that for all $t\in[0,\tau]$, it almost surely holds
$G(t|A,X)\geqslant\varepsilon.$
Figure 3 represents the causal graph when the study is randomized with conditionally independent censoring.
@@
Under dependent censoring, the Kaplan-Meier estimator as defined in Definition 2.1 can fail to estimate survival, in particular because Equation 5 no longer holds. Alternatives include plug-in G-formula estimators (Section 2.2.1) and unbiased transformations (Section 2.2.2).
2.2.1 The G-formula and the Cox Model
Because the censoring is now independent on the potential outcome conditionally to the covariates, it would seem natural to model the response of the survival time conditionally to these covariates too:
$\mu(x,a):=\mathbb{E}[T\wedge\tau|X=x,A=a].$
Building on Equation 4, one can express the RMST as a function of $\mu$:
@@ 2.2.1 The G-formula and the Cox Model
An estimator $\widehat{\mu}$ of $\mu$ would then straightforwardly yields an estimator of the RMST difference through the so-called G-formula plug-in estimator:
@@
We would like to stress that a G-formula approach works also in a observational context as the one introduced in Section 3.2. However, because the estimation strategies presented in the next sections rely on estimating nuisance parameters, and that this latter task is often done in the same way as we estimate the conditional response $\mu$, we decided to not delay the introduction of the G-formula any further, and we present below a few estimation methods for $\mu$. These methods are subdivided in two categories: $T$-learners (i.e. two learners), where $\mu(\cdot,1)$ is estimated separately from $\mu(\cdot,0)$, and $S$-learner (i.e. single learner), where $\widehat{\mu}$ is obtained by fitting a single model based on covariates $(X,A)$.
Cox Model. There are many ways to model $\mu$ in a survival context, the most notorious of which being the Cox proportional hazards model (Cox,, 1972). It relies on a semi-parametric model of the conditional hazard functions and specifies the function under treatment $a$, $\lambda^{(a)}(t|X)$, as
@@
where $\lambda^{(a)}_{0}$ is a baseline hazard function under treatment $a$ and $\beta^{(a)}$ has the same dimension as the vector of covariate $X$ (We use the superscript $(a)$ here to emphasize that a separate Cox model is fit for each treatment level, since the treatment assignment $A$ is not included as a covariate in $X$). The conditional survival function then takes the simple form (in the continuous case)
$S^{(a)}(t|X)=S^{(a)}_{0}(t)^{\exp(X^{\top}\beta^{(a)})},$
where $S^{(a)}_{0}(t)$ is the survival function associated with $\lambda_{0}^{(a)}$. The vector $\beta^{(a)}$ is classically estimated by maximizing the so-called partial likelihood function as introduced in the original paper of Cox, (1972). The cumulative baseline hazard function can be estimated through the Breslow’s estimator (Breslow,, 1974):
$\widehat{\Lambda}_{0}^{(a)}(t)=\sum_{\Delta_{i}=1,\widetilde{T}_{i}\leqslant t}\frac{1}{\displaystyle\sum_{\widetilde{T}_{j}\geqslant\widetilde{T}_{i}}\exp(X_{j}^{\top}\widehat{\beta}^{(a)})}$
where $\widehat{\beta}^{(a)}$ is a partial likelihood maximizer. One can show that $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ is the MLE of the true likelihood, when $\widehat{\Lambda}_{0}^{(a)}$ is optimized over all step functions of the form
$\Lambda_{0}(t):=\sum_{\Delta_{i}=1}h_{i},\quad h_{i}\in\mathbb{R}^{+}.$
This fact was already hinted in the original paper by Cox and made rigorous in many subsequent papers, see for instance Fan et al., (2010). Furthermore, if the true distribution follows a Cox model, then both $\widehat{\beta}^{(a)}$ and $\widehat{\Lambda}_{0}^{(a)}$ are strongly consistent and asymptotically normal estimator of the true parameters $\beta^{(a)}$ and $\Lambda^{(a)}$, see Kalbfleisch and Prentice, (2002), Sec 5.7. When using a $T$-learner approach, one simply finds $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ for $a\in\{0,1\}$ by considering the control group and the treated group separately. When using a $S$-learner approach, the treatment status $A$ becomes a covariate a the model becomes
$\lambda(t|X,A)=\lambda_{0}(t)\exp(X^{\top}\beta+\alpha A).$} (11)
for some $\alpha\in\mathbb{R}$. One main advantage of Cox model is that it makes it very easy to interpret the effect of a covariate on the survival time. If indeed $\alpha>0$, then the treatment has a negative effect on the survival times. Likewise, if $\beta_{i}>0$, then the $i$-th coordinate of $X$ has a negative effect as well. We finally mention that the hazard ratio takes a particularly simple form under the later model since
$\theta_{\mathrm{HR}}=e^{\alpha}.$
In particular, it does not depends on the time horizon $\tau$, and is thus sometimes referred to as proportional hazard. Figure 4 illustrates the estimation of the difference in Restricted Mean Survival Time using G-formula with Cox models.
Finally, while we have focused here on Cox-based specifications due to its widespread use in biomedical research. However, alternative approaches exist. For example, Accelerated Failure Time (AFT) models (Saikia and Barman,, 2017) provide a semi-parametric alternative that directly models the survival time and does not rely on the proportional hazards (PH) assumption.
Figure 4: Illustration of the G-formula for estimating $\theta_{\mathrm{RMST}}$ in an RCT when only one covariate $X_{1}$ influences the outcome.
Weibull Model. A popular parametric model for survival is the Weibull Model, which amounts to assume that
$\lambda^{(a)}(t|X)=\lambda_{0}^{(a)}(t)\exp(X^{\top}\beta)$
where $\lambda_{0}^{(a)}(t)$ is the instantaneous hazard function of a Weibull distribution, that is to say a function proportional to $t^{\gamma}$ for some shape parameter $\gamma>0$. We refer to Zhang, (2016) for a study of this model.
Survival Forests. On the non-parametric front, we mention the existence of survival forests (Ishwaran et al.,, 2008).
2.2.2 Censoring-unbiased transformations
@@
for two wisely chosen transformations $\phi_{0}$ and $\phi_{1}$, and where
$\Delta^{\tau}:=\mathbb{I}\{T\wedge\tau\leqslant C\}=\Delta+(1-\Delta)\mathbb{I}(\widetilde{T}\geqslant\tau)$} (13)
is the indicator of the event where the individual is either uncensored or censored after time $\tau$. The idea behind the introduction of $\Delta^{\tau}$ is that because we are only interested in computing the expectation of the survival time thresholded by $\tau$, any censored observation coming after time $\tau$ can in fact be considered as uncensored ($\Delta^{\tau}=1$).
A censoring-unbiased transformation $T^{*}$ shall satisfy: for $a\in\{0,1\}$, it holds
@@
A notable advantage of this approach is that it enables the use of the full transformed dataset $(X_{i},A_{i},T^{*}_{i})$ as if no censoring occurred. Because it holds
@@
One can derive an estimator of the RMST difference from any censoring-unbiased transformation $T^{*}$ as:
$\widehat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{i}$} (16)
@@
The following proposition establishes the theoretical guarantees for the class of estimators constructed from bounded censoring-unbiased transformations:
Proposition 2.1.
@@
It is natural in a RCT setting to assume that probability of being treated $\pi$ is known. If not, it is usual to replace $\pi$ by its empirical counterpart $\widehat{\pi}=n_{1}/n$ where $n_{a}=\sum_{i}\mathds{1}\{A=a\}$. The resulting estimator takes the form
@@
While this estimator introduces a slight bias due to the estimation of $\pi$ (see for instance Colnet et al., (2022), Lemma 2), it remains strongly consistent and asymptotically normal under the same assumptions:
Proposition 2.2.
@@
The two most popular transformations are Inverse-Probability-of-Censoring Weighting (Koul et al.,, 1981) and Buckley-James (Buckley and James,, 1979), both illustrated in Figure 5 and detailed below. In the former, only non-censored observations are considered and they are weighted while in the latter, censored observations are imputed with an estimated survival time.
@@
The Inverse-Probability-of-Censoring Weighted (IPCW) transformation, introduced by Koul et al., (1981) in the context of censored linear regression, involves discarding censored observations and applying weights to uncensored data. More precisely, we let
@@
where we recall that $G(t|X,A):=\mathbb{P}(C\geqslant t|X,A)$ is the left limit of the conditional survival function of the censoring. This estimator assigns higher weights to uncensored subjects within a covariate group that is highly prone to censoring, thereby correcting for conditionally independent censoring and reducing selection bias (Howe et al.,, 2016).
We establish in Appendix 9, Section 9.2, (Proposition 9.3) that the IPCW transformation is indeed a censoring-unbiased transformation in the sense of Equation 14.
The IPCW depends on the unknown conditional survival function of the censoring $G(\cdot|X,A)$, which thus needs to be estimated. Estimating conditional censoring or the conditional survival function can be approached similarly, as both involve estimating a time—whether for survival or censoring. Consequently, we can use semi-parametric methods, such as the Cox model, or non-parametric approaches like survival forests, and we refer to Section 2.2.1 for a development on these approaches. Once an estimator $\widehat{G}(\cdot|A,X)$ of the later is provided, one can construct an estimator of the difference in RMST based on Equation 16 or Equation 17
$\widehat{\theta}_{\mathrm{IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{IPCW},i},$} (19)
or
@@
where we recall that $n_{a}:=\#\{i\in[n]~|~A_{i}=a\}$. By Proposition 2.1, Proposition 2.2 and Proposition 9.3, we easily deduce that $\widehat{\theta}_{\mathrm{IPCW}}$ is asymptotically consistent as soon as $\widehat{G}$ is.
@@
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{\theta}_{\mathrm{IPCW}}$, either as in defined in Equation 19 or in Equation 20.
This result simply comes from the fact that $\widehat{\theta}_{\mathrm{IPCW}}$ depends continuously on $\widehat{G}$ and that $G$ is lower-bounded (Assumption 2.6 .Conditionally independent censoring). Surprisingly, we found limited use of this estimator in the literature, beside its first introduction in Koul et al., (1981). This could potentially be explained by the fact that, empirically, we observed that this estimator is highly variable. Consequently, we do not explore its properties further and will not include it in the numerical experiments. A related and more popular estimator is the IPCW-Kaplan-Meier, defined as follows.
@@
$D_{k}^{\mathrm{IPCW}}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
We show in Proposition 9.4 (Appendix 9) that the oracle version of this estimator is strongly consistent and asymptotically normal. These properties also hold for the estimator using a consistently estimated $\widehat{G}$.
@@
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{S}_{\mathrm{IPCW}}(t|A=a)$.
Because the evaluation of $N_{k}^{\textrm{IPCW}}(a)$ now depends on information gathered after time $t_{k}$ (through the computation of the weights), the previous properties on the absence of bias (Proposition 9.1) and on the derivation of the asymptotic variance (Proposition 9.2) do not hold in this setting. Whether its bias is exponentially small and whether the asymptotic variance can be derived in a closed form are questions left open. We are also not aware of any estimation schemes for the asymptotic variance in this context. In the case where we do not have access to the oracle survival function $G$, we can again still achieve consistency if the estimator $\widehat{G}(\cdot|A,X)$ that we provide is consistent.
@@
One weakness of the IPCW transformation is that it discards all censored data. The Buckley-James (BJ) transformation, introduced by Buckley and James, (1979), takes a different path by leaving all uncensored values untouched, while replacing the censored ones by an extrapolated value. Formally, it is defined as follows:
@@
where $S(t|X,A=a):=\mathbb{P}(T^{(a)}>t|X)$ is the conditional survival function. We refer again to Figure 5 for a diagram of this transformation.
We show in Appendix 9 (Proposition 9.5) that this transformation satisfies the censoring-unbiasedness condition defined in Equation 14.
Again, the BJ transformation depends on a nuisance parameter (here $Q_{S}(\cdot|X,A)$) that needs to be estimated. We again refer to Section 2.2.1 for a brief overview of possible estimation strategies for $Q_{S}$. Once provided with an estimator $\widehat{Q}_{S}(\cdot|A,X)$, a very natural estimator of the RMST based on the BJ transformation and on Equation 16 or Equation 17 would be
$\widehat{\theta}_{\mathrm{BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{BJ},i},$} (23)
or
@@
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{Q}_{S}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{\theta}_{\mathrm{BJ}}$ defined as in Equation 23 or Equation 24.
The proof is again a mere application of Propositions 2.1, 2.2 and 9.5, and relies on the continuity of $S\mapsto Q_{S}$.
Additionally, the Buckley-James transformation enjoys a notable optimality property: it minimizes the mean squared error among all square-integrable censoring-unbiased transformations. We restate formally as Theorem 9.1 in Appendix 9. This result, originally stated without proof in Fan and Gijbels, (1994), is proved there for completeness.

### v3

@@ 2.2 Conditionally independent censoring
Conditionally independent censoring relaxes the assumption of independent censoring by allowing censoring to depend on treatment or covariates — for instance, when treated patients withdraw from a study due to side effects.
@@
(Conditionally Independent Censoring) $C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X,A$
@@
(Conditional Positivity / Overlap for Censoring) There exists $\varepsilon>0$ such that for all $t\in[0,\tau]$, it almost surely holds $G(t|A,X)\geqslant\varepsilon.$
Figure 3 represents the causal graph when treatment assignment is randomized and when there is a conditionally independent censoring mechanism.
@@
Under dependent censoring, the Kaplan-Meier estimator as defined in Definition 2.1 is not suited to estimate survival. Alternatives include plug-in G-formula estimators (Section 2.2.1) and unbiased transformations (Section 2.2.2).
2.2.1 The G-formula
Building on Equation 2, one can express the RMST conditioning on the covariates:
@@ 2.2.1 The G-formula
with $\mu(x,a):=\mathbb{E}[T\wedge\tau|X=x,A=a].$
An estimator of the RMST difference via the G-formula arises by plugging in any estimator $\widehat{\mu}$ of $\mu$:
@@
The G-formula accommodates right censoring under conditionally independent censoring, without specifying a separate censoring model: the target (e.g., $S^{(a)}(t)$ or $\mathrm{RMST}^{(a)}$) is expressed as a functional of the conditional outcome regression $\mu(\cdot)$. The same G-formula can be applied to observational settings (in Section 3.2) to address non-random treatment assignment. Different strategies can be used to estimate $\mu$, ranging from semi-parametric models such as the Cox model to fully non-parametric approaches like survival forests. For any chosen model, implementation follows one of two paradigms: T-learner approaches (i.e., two separate learners), where $\mu(\cdot,1)$ and $\mu(\cdot,0)$ are estimated separately; and S-learner approaches (i.e., a single learner), where $\widehat{\mu}$ is obtained by fitting one model using covariates $(X,A)$.
Cox Model. The Cox proportional hazards model (Cox1972) relies on a semi-parametric model of the conditional hazard functions and specifies the function under treatment $a$, $\lambda^{(a)}(t|X)$, as
@@
where $\lambda^{(a)}_{0}$ is a baseline hazard function under treatment $a$ and $\beta^{(a)}$ has the same dimension as the vector of covariate $X$. The model is semi-parametric in the sense that the baseline hazard is left unspecified, while the effect of covariates enters through a multiplicative proportional hazards structure. Alternative semi-parametric models include Accelerated Failure Time (AFT) (saikia2017) which directly models the survival time and does not rely on the proportional hazards (PH) assumption.
Weibull Model. A popular parametric model for survival is the Weibull Model, which amounts to assume that $\lambda^{(a)}(t|X)=t^{\gamma(a)}\exp(X^{\top}\beta)$ where $\gamma(a)>0$. We refer to zhang2016 for a study of this model.
Survival Forests. On the nonparametric side, survival forests (ishwaran2008) aggregate survival trees that use log-rank splitting and terminal-node Nelson–Aalen estimates to learn $S(t\mid X)$ under right censoring, capturing nonlinearities and high-order interactions without parametric form.
2.2.2 Censoring-unbiased transformations (CUT)
@@
for two wisely chosen transformations $\phi_{0}$ and $\phi_{1}$, and where $\Delta^{\tau}:=\mathbb{I}\{T\wedge\tau\leqslant C\}=\Delta+(1-\Delta)\mathbb{I}(\widetilde{T}\geqslant\tau)$ indicates whether an individual is either uncensored or censored after time $\tau$. Since we only need the expectation of survival time thresholded at $\tau$, any censoring beyond $\tau$ can be treated as uncensored ($\Delta^{\tau}=1$). A censoring-unbiased transformation $T^{*}$ is defined such that, for any $a\in\{0,1\}$,
@@
A notable advantage of this approach is that it enables the use of the full transformed dataset $(X_{i},A_{i},T^{*}_{i})$ as if no censoring occurred. Moreover, by the reweighting identity
@@
this equality links the arm-specific mean of $T^{*}$ to its full-sample inverse-probability–weighted representation. The RMST difference can be estimated by
$\widehat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{i},$} (8)
@@
The following proposition establishes the theoretical guarantees for the class of estimators constructed from bounded censoring-unbiased transformations.
Proposition 2.1 (Consistency of CUT as RMST estimator with known $\pi$).
@@
It is natural in a RCT setting to assume that probability of being treated $\pi$ is known. If not, it is usual to replace $\pi$ by its empirical counterpart $\widehat{\pi}=n_{1}/n$ where $n_{a}=\sum_{i}\mathbbm{1}\{A=a\}$. The resulting estimator takes the form
@@
While this estimator introduces a slight bias due to the estimation of $\pi$ (see for instance colnet2022, Lemma 2), it remains strongly consistent and asymptotically normal under the same assumptions:
Proposition 2.2 (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).
@@
For the later, we consider only estimator with estimated probability of being treated in Equation 9.
The two most popular transformations are Inverse-Probability-of-Censoring Weighting (koul1981) and Buckley-James (buckley1979), both illustrated in Figure 4 and detailed below.
@@
In the former, only non-censored observations are considered and they are weighted while in the latter, censored observations are imputed with an estimated survival time.
@@
The Inverse-Probability-of-Censoring Weighted (IPCW) transformation, introduced by koul1981 in the context of censored linear regression, involves discarding censored observations and applying weights to uncensored data (see Figure 4). More precisely, we let
@@
where we recall that $G(t|X,A):=\mathbb{P}(C\geqslant t|X,A)$ is the left limit of the conditional survival function of the censoring. This estimator assigns higher weights to uncensored subjects within a covariate group that is highly prone to censoring, thereby correcting for conditionally independent censoring and removing selection bias (howe2016).
Using proposition 2.1 and 2.2, we show in Appendix that the IPCW transformation is a censoring-unbiased transformation in the sense of Equation 6 (Proposition 6.3).
The IPCW depends on the unknown conditional survival function of the censoring $G(\cdot|X,A)$, which thus needs to be estimated. Estimating conditional censoring or the conditional survival function can be approached similarly, as both involve estimating a time, whether for survival or censoring. Consequently, we can use parametric methods, or semi-parametric methods, such as the Cox model, or non-parametric approaches like survival forests (see Section 2.2.1). Once $\widehat{G}(\cdot|A,X)$ is provided, one can construct an estimator of the difference of RMST based on Equation 9,
@@
Alternatively, instead of rescaling by $n_{a}$, the estimator can be stabilized by dividing by the sum of the weights, yielding a Hájek-type normalization that reduces variance without affecting consistency (Hajek1971; Robins2000; Hernan2020).
By Proposition 2.1 ( (Consistency of CUT as RMST estimator with known $\pi$).), Proposition 2.2 ( (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).) and Proposition 6.3 ( (IPCW is a Censoring Unbiased Transformation).) (in Appendix A.1.2. Proofs of Section 2.2), we easily deduce that $\widehat{\theta}_{\mathrm{IPCW}}$ is asymptotically consistent as soon as $\widehat{G}$ is.
@@
This result simply comes from the fact that $\widehat{\theta}_{\mathrm{IPCW}}$ depends continuously on $\widehat{G}$ and that $G$ is lower-bounded (Assumption 2.6 .Conditionally independent censoring). Surprisingly, we found limited use of this estimator in the literature, beside its first introduction in koul1981.
A more common variant is the IPCW–Kaplan–Meier (IPCWrobins; IPCW_consistency; IPCW), defined as follows. In this approach, censoring weights are dynamically evaluated at each distinct event time, allowing the estimator to update risk sets and weighted failures throughout the follow-up period.
@@
$D_{k}^{\mathrm{IPCW}}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\frac{1}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
We show in Proposition 6.4 (Appendix A.1. Appendix A: Proofs) that the oracle version of this estimator is strongly consistent and asymptotically normal. These properties also hold for the estimator using a consistently estimated $\widehat{G}$.
@@
Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{S}_{\mathrm{IPCW}}(t|A=a)$.
Because the evaluation of $N_{k}^{\textrm{IPCW}}(a)$ now depends on information gathered after time $t_{k}$, the previous derivations of bias (Appendix A.1.1. Proofs of Section 2.1, proposition 6.1) and asymptotic variance (Appendix A.1.1. Proofs of Section 2.1, proposition 6.2) no longer hold. Overgaard2025 derived asymptotic variance expressions for related Kaplan–Meier–based inverse probability-of-censoring weighted estimators using influence-function representations. While these results theoretically extend to our IPCW–Kaplan–Meier formulation, the resulting expressions remain difficult to interpret or implement in practice. Based on our own results, we can nevertheless establish that when $G$ is unknown, consistency still holds provided that the estimator $\widehat{G}(\cdot|A,X)$ is itself consistent.
@@
The Buckley-James (BJ) transformation, introduced by buckley1979, takes a different path by leaving all uncensored values untouched, while replacing the censored ones by an extrapolated value (see Figure 4). Formally, it is defined as follows:
@@
and $S(t|X,A=a):=\mathbb{P}(T^{(a)}>t|X)$ is the conditional survival function.
This transformation satisfies the censoring-unbiasedness condition defined in Equation 6 (Proposition 6.5 in Appendix A.1. Appendix A: Proofs). The BJ transformation depends on a nuisance parameter ($Q_{S}(\cdot|X,A)$) that needs to be estimated. Estimation strategies described in Section 2.2.1 can be used. Once provided with an estimator $\widehat{Q}_{S}(\cdot|A,X)$, a natural estimator of the RMST based on the BJ transformation and Equation 9 is
@@
Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{Q}_{S}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{\theta}_{\mathrm{BJ}}$ defined as in Equation 14.
The proof is an application of Propositions 2.1 ( (Consistency of CUT as RMST estimator with known $\pi$).), 2.2 ( (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).) and 6.5 ( (BJ is a censoring unbiased transformation). in Appendix, Section A.1.2. Proofs of Section 2.2), and relies on the continuity of $S\mapsto Q_{S}$. Additionally, the Buckley-James transformation enjoys a notable optimality property.
Theorem 2.1 (BJ minimizes the mean squared error among CUT).
For any transformation $T^{*}$ of the form 5, it holds
$\mathbb{E}[(T^{*}_{\mathrm{BJ}}-T\wedge\tau)^{2}]\leqslant\mathbb{E}[(T^{*}-T\wedge\tau)^{2}].$
This result, originally stated without proof in fan1994, is proved there for completeness in Appendix, Section A.1.2. Proofs of Section 2.2.